A=pir^2
for pi=22/7 and r=7/4

Answer by Cromlix(4381)

Area = Pi x r^2
Area = 22/7 x (7/4)^2
Area = 22/7 x 49/16
Area = 1078/112
Area = 9 5/8
Hope this helps:-)